=== 2019/4/17(Wed.) ===

==== Schedule ====
 * Center
  * --(3/25: Venting)--
  * --(3/27-4/9: Fixing PR2 or others)--
   * --(4/4 -4/5: Gluing)--
   * --(4/8: alignment check)--
   * --(4/8-10: OSEM adjustment)--
  * --(4/10: alignment check)--
  * --(4/11: VIS final health check)--
  * --(4/11-12: Closing)--
  * --(4/12: Pump down)--
  * 4/29,30: Leak check

 * ITMX: 
  * --(3/19: Cooling stop)--
  * --(3/25: Venting start)--
  * --(3/27: Opning chamber)--
  * --(3/27~: Fixing thermal touch)--
  * --(4/2~: closing chamber and start pumping down)--
  * --(4/10: Cooling starts.)--

 * Xend
  * --(3/25: closing and pump down)--
  * --(4/8: Leak check)--
  * --(4/9-10: venting)--
  * --(4/11: Exchange a broken LED for Photo sensor)--
  * --(4/12: closing chamber and start pumping down)--
  * 4/22: Leak check
  * 4/22: Cooling start

 * Yend
  * --(3/26: Venting start)--
  * --(3/28~: Fixing short inside chamber)--
  * --(4/3: Fixing F0 actuator)--
  * --(4/4~: closing chamber and start pumping down.)--
  * --(4/10: Cooling starts.)--

 * April: Measurement of TF, payload diagonalization on ITMX
 * End of April~: DRMI commissioning
 * End of April~: Y arm commissioning
 * ~End of May: FPMI commissioning

==== Note ====
 * No 1pm Commissioning meeting this week(4/15-19).

==== Night work ====
 * Alignment (Nakano, Enomoto, Sugimoto, Yokozawa)

==== Weekend ====

==== Report of work ====
 * Type Bp: Based on measurements of TF, PR2 and PR3 looks healthy. PRM needs more time.
 * Type A: ETMX may have swapped cable.
 * SR3 TF seems to be healthy.
 * CRY: ITMY cannot show role resonance. It may have touching. No need to open chamber again.

 * MIF: both end gate valves were opened.
  * GR reached ETMX.
  * GR reached ETMY and confirmed hitting center.

==== Work plan ====

 * Health check Payload of ETMX will be performed.
 * CAL: maintenance ETMX circuit today and tomorrow.